OpenStudy (yrivers36):

Please help me to understand. y ^(-1/5) I changed problem to 1/y^(1/5) after this point I am not sure how to solve once its in radical form

OpenStudy (amistre64):

its an expression, not an equation

OpenStudy (amistre64):

what constitutes a solution to a statement?

OpenStudy (savvy):

you cannot solve it any further until and unless it is equal to something...

OpenStudy (anonymous):

guessing y^(1/5)=\[\sqrt[5]{y}\]

OpenStudy (yrivers36):

Ok so I did have it right. Thank you so much

OpenStudy (anonymous):

I just changed the form of the equation. hint is look at the fraction and that what the sqrt will be powered to

OpenStudy (yrivers36):

ok so then it will be the denominator and one the numerator

OpenStudy (anonymous):

for the most part yes.

OpenStudy (yrivers36):

Thank you for your help

OpenStudy (anonymous):

y^(1/2)= is just \[\sqrt{y}\] because the sqrt is equal to second power that's why if you ever want to get rid of the sqrt just sq it

OpenStudy (yrivers36):

it was y^(-1/5)

OpenStudy (anonymous):

i know. i was just giving an example

OpenStudy (yrivers36):

oh ok sorry

OpenStudy (anonymous):

no worries have a good one

OpenStudy (yrivers36):

you too